YZ Corporation acquired in 1/1/2010 a second-hand piece of machinery by EF Corporation in cash. In the records of the EF Corporation the book value of the machinery at the date of the sale was €15,000. The invoice price of the machinery was €14,000, while the EF corporation offered a cash discount of €1,000. In addition the EF Corporation paid €150 and €470 for transportation expenses and installation costs respectively. The piece of machinery was ready for its intended use on 1/1/2010. According to its technical specification the machinery would have been fit for use for 20 years. The management of XYZ corporation intended to use the machinery until 31/12/2019. The residual value of the machinery was estimated to €1,200. In 1/1/2018 a major overhaul of the machinery took place. As a consequence of the overhaul the production capacity of the machinery increased and the management of XYZ decided that the machinery would have been used until 31/12/2022 when the residual value of the machinery was estimated to be insignificant. The cost of the overhaul was €1,500. In 1/1/2020 the machinery suffered a major break-down and the management of XYZ decided to retire the particular machinery at the same date. XYZ Corporation employed the straight-line method for the calculation of the annual depreciation charges. The firm’s accounting period commences at 1st January and ends at 31st December of each year.

Determine the accounting entry that XYZ should make in order to record the depreciation charge for fiscal period 2018:

Τhe correct answer is Debit: Depreciation 1,012, Credit: Machinery-accumulated depreciation 1,012 (Not 1180)

but why?
Answer #1

Acquisition cost of machinery = €14000 – €1000 = €13000

Annual straight-line depreciation = (€13000 – €1200)/10 years = €1180

Depreciation charged from 1/1/2010 to 31/12/2017 = €1180 x 8 = €9440

Book value on 1/1/2018 = €13000 – €9440 = €3560

Depreciable cost for 2018 = €3560 + €1500 = €5060

Remaining useful life from 1/1/2018 to 31/12/2022 = 5 years

Depreciation for 2018 = €5060/5 years = €1012

The depreciation is computed prospectively over the revised remaining useful life of the machinery.